# VAMO: Efficient Large-Scale Nonconvex Optimization via Adaptive Zeroth Order Variance Reduction

### v1

@@ 7 Conclusion
In this paper, we propose a hybrid FO and ZO variance-reduced algorithm, VAMO, for nonconvex optimization. We demonstrate that compared to FO-SGD, our algorithm improves the convergence rate from $O(1/\sqrt{T})$ to a linear rate of $O(1/T)$, achieving convergence performance similar to FO-SVRG. Compared to ZO algorithms, our method maintains convergence performance independent of the problem dimension $d$, making it effective for optimizing high-dimensional problems. However, due to the use of two-point ZO gradient estimation, our convergence result includes an additional error term $O(1/b)$. To mitigate this, we introduce a multi-point ZO gradient estimation variant, which reduces this error. Unlike previous purely FO or ZO methods, our hybrid approach leverages the advantages of both, enabling a more flexible trade-off between computational efficiency and convergence performance. This makes it more adaptable to real-world applications with complex constraints. Our theoretical analysis and empirical evaluations, including comparisons with state-of-the-art methods, demonstrate the effectiveness of our approach.
Appendix A Appendix / supplemental material
A.1 ZO gradient estimator
@@
    where ^$\hat{\nabla}f(\mathbf{x})$ is the multi-point gradient estimate.
@@
See the proof of Lemma 2 in liu2018zeroth
A.2 Second-Order Moment of the Hybrid Gradient Estimator
@@
where the first inequality holds due to Lemma 4. Based on (26), we note that the following holds
@@
Based on (29) and applying Lemma 1 and Lemma 3, the first term at the right hand side (RHS) of (28) yields
@@
Substituting (30) into (28), we obtain
@@
Similar to Lemma 1, we introduce a smoothing function $f_{i,\mu}$ of $f_{i}$, and continue to bound the second term at the right hand side (RHS) of (32). This yields
@@
$\leq 12d\mathbb{E}[\|\nabla f_{i}(\mathbf{x}_{k}^{s})\|_{2}^{2}]+(12d+3)L^{2}\mathbb{E}\left[\|\mathbf{x}_{0}^{s}-\mathbf{x}_{k}^{s}\|_{2}^{2}\right]+\frac{9}{4}L^{2}d^{2}\mu^{2}$
$\leq 24d\mathbb{E}\left[\|\nabla f_{i}(\mathbf{x}_{k}^{s})-\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+24d\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]$
$+(12d+3)L^{2}\mathbb{E}\left[\|\mathbf{x}_{0}^{s}-\mathbf{x}_{k}^{s}\|_{2}^{2}\right]+\frac{9}{4}L^{2}d^{2}\mu^{2}$
$\leq 24d\sigma^{2}+24d\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+(12d+3)L^{2}\mathbb{E}\left[\|\mathbf{x}_{0}^{s}-\mathbf{x}_{k}^{s}\|_{2}^{2}\right]+\frac{9}{4}L^{2}d^{2}\mu^{2},$
@@
We bound the first term at the right hand side (RHS) of (32). This yields
@@ A.3 Proof of Theorem 1
where the last equality holds due to $x_{k+1}^{s}=x_{k}^{s}-\eta_{k}v_{k}^{s}$. Since $x_{k}^{s}$ and $x_{0}^{s}$ are independent of $\mathcal{I}$ and random directions $u$ used for ZO gradient estimates, from (18) we obtain
@@
Combining (38) and (39), we have
@@
At RHS of (40), the upper bound on $\mathbb{E}\left[\|{\mathbf{v}}_{k}^{s}\|_{2}^{2}\right]$ is given by Proposition 1,
@@
In (40), we further bound $\mathbb{E}\left[\|\mathbf{x}_{k+1}^{s}-\mathbf{x}_{0}^{s}\|_{2}^{2}\right]$ as,
@@
for some $c_{k}>0$, Substituting (40) and (42) into $R_{k+1}^{s}$, we obtain
@@
Moreover, substituting (41) into (44), we have
$R_{k+1}^{s}\leq$ $\mathbb{E}\left[f(\mathbf{x}_{k}^{s})\right]\boldsymbol{-}\left(\eta_{k}-\frac{c_{k\boldsymbol{+}1}\eta_{k}}{\beta_{k}}\right)\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+\left(c_{k\boldsymbol{+}1}+c_{k\boldsymbol{+}1}\beta_{k}\eta_{k}\right)\mathbb{E}\left[\|\mathbf{x}_{k}^{s}-\mathbf{x}_{0}^{s}\|_{2}^{2}\right]$ (45)
@@
Based on the definition of $c_{k}=c_{k+1}+\beta_{k}\eta_{k}c_{k+1}+\left(\frac{L}{2}\eta_{k}^{2}+c_{k+1}\eta_{k}^{2}\right)\frac{12(4d+1)L^{2}\delta_{n}}{b}\alpha^{2}$ and the definition of $R_{k}^{s}$ in (43), we can simplify the inequality (45) as
@@
Taking a telescopic sum for (47), we obtain
@@
where $\chi_{m}=\sum_{k=0}^{m-1}\chi_{k}$. It is known from (43) that,
@@
where the last equality used the fact that $c_{m}=0$, since $\bar{\mathbf{x}}_{s\boldsymbol{-}1}=\mathbf{x}_{0}^{s}$ and $\bar{\mathbf{x}}_{s}=\mathbf{x}_{m}^{s}$, we obtain
@@ A.4 Proof of Corollary 1
where $\theta=\beta\eta+\frac{12(1+4d)L^{2}\delta_{n}\eta^{2}}{b}\alpha^{2}$. The recursive formula (53) implies that $c_{k}\leq c_{0}$ for any $k$, and
@@
where we have used the fact that $\delta_{n}\leq 1$, Substituting (55) into (54), we have
@@
We recall from (8) and (9) that
@@
From (56) and the definition of $\beta$, we have
@@
Substituting (LABEL:eq:gamma_first_bound) and (60) into (58), we obtain
@@
Next, we find the upper bound on $\chi_{m}$ in (9) given the current parameter setting and $c_{k}\leq c_{0}$,
@@
where in the big $O$ notation, we only keep the dominant terms and ignore the constant numbers that are independent of $d$, $b$, and $T$
Substituting (62) and (65) into (7), we have
@@ A.5 Proof of Theorem 2
Following together with (24), we can obtain that
@@
Substituting (68) and (36) into (67), we have:
@@
Substituting (69) into (44), we have:
$R_{k+1}^{s}\leq$ $\mathbb{E}\left[f(\mathbf{x}_{k}^{s})\right]\boldsymbol{-}\left(\eta_{k}-\frac{c_{k\boldsymbol{+}1}\eta_{k}}{\beta_{k}}\right)\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+\left(c_{k\boldsymbol{+}1}+c_{k\boldsymbol{+}1}\beta_{k}\eta_{k}\right)\mathbb{E}\left[\|\mathbf{x}_{k}^{s}-\mathbf{x}_{0}^{s}\|_{2}^{2}\right]$ (70)
@@
Based on the definition of $c_{k}=\left(1+\beta_{k}\eta_{k}\right)c_{k\boldsymbol{+}1}+\left(\frac{L}{2}+c_{k+1}\right)\left(1+\frac{4d}{q}\right)\frac{12L^{2}\delta_{n}\eta_{k}^{2}\alpha^{2}}{b}$ and $R_{k}^{s}$ given by (43), we can simplify (70) to
$R_{k+1}^{s}$ $\leq R_{k}^{s}\boldsymbol{-}\left(\eta_{k}-\frac{c_{k\boldsymbol{+}1}\eta_{k}}{\beta_{k}}\right)\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]$ (71)
@@
Based on (71) and the following argument in, we can achieve
@@
$\theta=\rho+\left(q+4d\right)\frac{12\delta_{n}q\rho^{2}}{bd^{2}}\leq\rho+12\rho\left(\frac{q^{2}}{d^{2}}+4\frac{q}{d}\right)\leq\rho+\frac{60\rho q}{d}$ (74)
Substituting (74) into (73), we have:
@@
The second inequality holds if we let $\rho\leq\frac{1}{120}$ Substituting (78) and (77) into (76), we can get
@@
Substituting (79) and (82) into (7), we have

### v2

@@ 7 Conclusion
In this paper, we propose a hybrid FO and ZO variance-reduced algorithm, VAMO, for nonconvex optimization. We demonstrate that compared to FO-SGD, our algorithm improves the convergence rate from $O(1/\sqrt{T})$ to a linear rate of $O(1/T)$, achieving convergence performance similar to FO-SVRG. Compared to ZO algorithms, our method maintains convergence performance independent of the parameter dimension $d$, making it effective for optimizing high-dimensional problems. However, due to the use of two-point ZO gradient estimation, our convergence result includes an additional error term $O(1/b)$. To mitigate this, we introduce a multi-point ZO gradient estimation variant, which reduces this error. Unlike previous purely FO or ZO methods, our hybrid approach leverages the advantages of both, enabling a more flexible trade-off between efficiency and convergence performance. This makes it more adaptable to real-world applications with complex constraints. Our theoretical analysis and empirical evaluations, including comparisons with state-of-the-art methods, demonstrate the effectiveness of our approach.
Appendix A Detailed Discussion on Memory Efficiency
We have made a brief analysis of VAMO’s memory efficiency in Section 4.3. In this section, we analyse VAMO’s memory profile in detail and compare it against standard FO and ZO optimizers, theoretically and empirically.
A.1 Theoretical Analysis
We adapt the memory analysis framework of Zhang et al. (2024) to compare the peak memory consumption of different optimizers. Table 3 summarizes the decomposition into three components: the model (Weight Mem.), the optimizer states (Opt. State Mem.), and dynamic allocation for computing gradients and optimization (Dynamic Mem.).
Table 3: Comparison of the instant peak memory consumption of different optimizers when fine-tuning the full model. Here $|\mathbf{x}|$ denotes the memory required to store the model parameters in full precision, and $|\mathbf{x}_{l}|$ is that of the parameters in layer $l$. The symbol $|\mathbf{a}_{l}|$ denotes the memory of intermediate activations of one single sample at layer $l$, where $b$ and $n$ correspond to the mini-batch size and the total dataset size, respectively. In Dynamic Mem., $b\cdot|\mathbf{a}_{l}|$ represents the memory of saving mini-batch activations at layer $l$, while $|\mathbf{x}_{l}|$ also accounts for temporarily saved gradients, as they have the same size as the corresponding parameters (Zhang et al., 2024).
Optimizer Weight Mem. Dynamic Mem. (Grad.&Opt.) Opt. State Mem.
FO-SGD $|\mathbf{x}|$ $\sum_{l}\max\{b\cdot|\mathbf{a}_{l}|,|\mathbf{x}_{l}|\}$ 0
FO-SVRG (accumulation) $|\mathbf{x}|$ $\sum_{l}\max\{b\cdot|\mathbf{a}_{l}|,|\mathbf{x}_{l}|\}$ $2|\mathbf{x}|$
FO-SVRG (full batch at once) $|\mathbf{x}|$ $\sum_{l}\max\{n\cdot|\mathbf{a}_{l}|,|\mathbf{x}_{l}|\}$ $2|\mathbf{x}|$
FO-Adam $|\mathbf{x}|$ $\sum_{l}\max\{b\cdot|\mathbf{a}_{l}|,|\mathbf{x}_{l}|\}$ $3|\mathbf{x}|$
FO-Adagrad $|\mathbf{x}|$ $\sum_{l}\max\{b\cdot|\mathbf{a}_{l}|,|\mathbf{x}_{l}|\}$ $2|\mathbf{x}|$
VAMO $|\mathbf{x}|$ $\sum_{l}\max\{b\cdot|\mathbf{a}_{l}|,|\mathbf{x}_{l}|\}$ $|\mathbf{x}|$
ZO-SGD $|\mathbf{x}|$ $\max_{l}|\mathbf{x}_{l}|$ 0
The advantage of ZO methods is the removal of backpropagation, which eliminates the need to store intermediate activations $|\mathbf{a}_{l}|$} for each input example during the forward pass. In FO optimizers, the total activation memory grows as $b\cdot\sum_{l}|\mathbf{a}_{l}|$ for a mini-batch of size $b$, while ZO methods completely avoid this dependence on the batch size. This reduction in dynamic memory cost becomes especially significant when $b$ is large in large-scale settings. Moreover, ZO gradients can be estimated in a layer-wise manner, so only the gradients of the currently processed layer need to be stored, further lowering the peak memory footprint compared to FO methods (Zhang et al., 2024). For ZO-SGD, dynamic memory is dominated by a temporary copy of the parameters of a single layer for gradient estimation, resulting in a peak memory of $\max_{l}|\mathbf{x}_{l}|$, independent of batch size, and no extra optimizer state is required.
In contrast, all FO methods require storing intermediate activations or other temporary variables to compute gradients, which results in dynamic memory that grows with the mini-batch size. FO (Johnson & Zhang, 2013) additionally needs to compute the full gradient at the snapshot point. Although this full gradient can be obtained by accumulating gradients over smaller mini-batches, reducing dynamic memory of intermediate results per step to $b\cdot|\mathbf{a}_{l}|$, it comes at the expense of higher computational cost per epoch due to multiple passes over all mini-batches. Computing the full gradient in a single full batch would substantially increase dynamic memory of intermediate results to $n\cdot|\mathbf{a}_{l}|$, requiring a compute-memory trade-off as commonly done in FO-SVRG. In contrast, VAMO leverages ZO estimation to compute the full gradient without storing intermediate activations, so even when computing over the full batch at once, the peak dynamic memory is only $\max_{l}|\mathbf{x}_{l}|$, the same as ZO-SGD, and no compute–memory trade-off is required. This makes VAMO highly memory-efficient for large mini-batches or large models.
Regarding optimizer state, adaptive methods such as Adam (Kingma & Ba, 2014) or Adagrad (Duchi et al., 2011) require additional memory for first- or second-order momentum, resulting in $2|\mathbf{x}|$$3|\mathbf{x}|$ extra memory per step. FO-SVRG also increases optimizer state by $2|\mathbf{x}|$ for storing the snapshot and full gradient (Johnson & Zhang, 2013). VAMO, on the other hand, only needs $|\mathbf{x}|$ for the full ZO gradient used in variance reduction, which is substantially smaller than Adam, Adagrad, or FO-SVRG. In memory-constrained scenarios, this makes VAMO a more favorable choice for fine-tuning large models.
A.2 Empirical Results Analysis
To empirically validate our theoretical analysis, we conducted full-parameter fine-tuning of a pre-trained RoBERTa-large model (Liu et al., 2019) and GPT-2 medium model (Radford et al., 2019) on the MNLI dataset (Williams et al., 2017) with FP32 (see Section 6). We measured the peak GPU memory consumption of VAMO, FO-SGD, FO-Adagrad and FO-Adam by varying the batch size (with sequence length fixed at 128). The detailed results are presented in Table 2 in Section 6.
We observe that VAMO’s peak memory consumption is consistently only marginally higher than FO-SGD, with the difference stemming from the overhead of storing the ZO full gradient for variance reduction. Importantly, as the batch size increases, the scaling trend of VAMO almost parallels that of FO-SGD, confirming that the intermediate results for computing mini-batch FO gradients dominates the memory cost in both methods. Notably, the ZO full gradient in VAMO does not require storing additional intermediate activations, further preventing the memory blow-up typically associated with FO-SVRG. In contrast, FO-Adam requires significantly more memory due to the additional storage of first- and second-moment estimates, resulting in an optimizer state of size $3|\mathbf{x}|$. This overhead makes FO-Adam substantially less memory efficient, especially for larger models, whereas VAMO maintains a footprint close to FO-SGD while avoiding the prohibitive costs of FO-SVRG. Taken together, these results highlight VAMO’s practical advantage: for a negligible increase in memory compared to FO-SGD, it achieves markedly better convergence, as shown in Section 6, while remaining far more memory efficient than FO-Adam. This balance makes VAMO particularly well suited for memory-constrained large-scale fine-tuning tasks.
Appendix B Detailed Discussing of Convergence and Complexity
In this section, we provide additional discussion of the convergence rates and computational complexities summarized in Table 1 (see Section 2). Table 1 summarizes the convergence rates and computational complexities of our proposed methods, referred to as VAMO and VAMO (multi-point) in the table alongside several FO and ZO algorithms.
For ZO methods, ZO-SVRG is listed with a complexity of $O(nS+bT)$ function queries. Among FO methods, FO-SGD has the lowest computational cost $O(bdT)$ but also exhibits the slowest convergence rate of $O(1/\sqrt{T})$. FO-SVRG improves convergence to $O(1/T)$ but increases the cost to $O(dnS+bdT)$ due to full gradient computations. Our proposed VAMO maintains a complexity of $O(nS+bdT)$, similar to ZO-SVRG in terms of $nS$ but replacing the $ndS$ full gradient cost of FO-SVRG with a cheaper $nS$ ZO estimation cost for snapshots, while achieving a fast $O(1/T+1/b)$ convergence rate. This makes its computational complexity significantly slower than FO-SVRG, especially when $d$ is large. The VAMO (multi-point) variant has a complexity of $O(qnS+bdT)$. Here, increasing $q$ (the number of ZO sampling directions) leads to higher complexity but also improves the convergence rate to $O(1/T+(1-q/d)^{2}/b)$, reducing the $O(1/b)$ error term and making its performance more comparable to FO-SVRG, particularly if $q\ll d$. This demonstrates that our proposed methods provide a flexible and often more efficient trade-off between computational cost and convergence performance compared to existing pure FO or ZO approaches. Our work further develops such an adaptive hybrid approach by specifically integrating ZO estimation within the SVRG structure, aiming to reduce the full-gradient cost while preserving strong convergence guarantees independent of dimensionality.
Appendix C ZO gradient estimator
@@
    where $\hat{\nabla}f(\mathbf{x})$ is the multi-point gradient estimate.
@@
See the proof of Lemma 2 in (Liu et al., 2018b)
Appendix D Second-Order Moment of the Hybrid Gradient Estimator
@@
where the first inequality holds due to Lemma 4. Based on equation 20, we note that the following holds
@@
Based on equation 23 and applying Lemma 1 and Lemma 3, the first term at the right hand side (RHS) of equation 22 yields
@@
Substituting equation 24 into equation 22, we obtain
@@
Similar to Lemma 1, we introduce a smoothing function $f_{i,\mu}$ of $f_{i}$, and continue to bound the second term at the right hand side (RHS) of equation 26. This yields
@@
[…]
$2]\displaystyle\leq 4d\mathbb{E}\left[\|\nabla f_{i}(\mathbf{x}_{k}^{s})-\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+4d\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right$]
[…]
@@
We bound the first term at the right hand side (RHS) of equation 26. This yields
@@ Appendix E Proof of Theorem 1
where the last equality holds due to $x_{k+1}^{s}=x_{k}^{s}-\eta_{k}v_{k}^{s}$. Since $x_{k}^{s}$ and $x_{0}^{s}$ are independent of $\mathcal{I}$ and random directions $u$ used for ZO gradient estimates, from equation 12 we obtain
@@
Combining equation 32 and equation 33, we have
@@
At RHS of equation 34, the upper bound on $\mathbb{E}\left[\|{\mathbf{v}}_{k}^{s}\|_{2}^{2}\right]$ is given by Proposition 1,
@@
In equation 34, we further bound $\mathbb{E}\left[\|\mathbf{x}_{k+1}^{s}-\mathbf{x}_{0}^{s}\|_{2}^{2}\right]$ as,
@@
for some $c_{k}>0$, Substituting equation 34 and equation 36 into $R_{k+1}^{s}$, we obtain
@@
Moreover, substituting equation 35 into equation 38, we have
$R_{k+1}^{s}\leq$ $\mathbb{E}\left[f(\mathbf{x}_{k}^{s})\right]\bm{-}\left(\eta_{k}-\frac{c_{k\bm{+}1}\eta_{k}}{\beta_{k}}\right)\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+\left(c_{k\bm{+}1}+c_{k\bm{+}1}\beta_{k}\eta_{k}\right)\mathbb{E}\left[\|\mathbf{x}_{k}^{s}-\mathbf{x}_{0}^{s}\|_{2}^{2}\right]$ (39)
@@
The definition of $c_{k}$ is given by
$c_{k}=c_{k+1}+\beta_{k}\eta_{k}c_{k+1}+\left(\frac{L}{2}\eta_{k}^{2}+c_{k+1}\eta_{k}^{2}\right)\frac{12(4d+1)L^{2}\delta_{n}}{b}\alpha^{2}$ (40)
Based on the definition of $c_{k}$ and the definition of $R_{k}^{s}$ in equation 37, we can simplify the inequality equation 39 as
@@
Taking a telescopic sum for equation 42, we obtain
@@
where $\chi_{m}=\sum_{k=0}^{m-1}\chi_{k}$. It is known from equation 37 that,
@@
where the last equality used the fact that $c_{m}=0$, since $\bar{\mathbf{x}}_{s\bm{-}1}=\mathbf{x}_{0}^{s}$ and $\bar{\mathbf{x}}_{s}=\mathbf{x}_{m}^{s}$, we obtain
@@ Appendix F Proof of Corollary 1
where $\theta=\beta\eta+\frac{12(1+4d)L^{2}\delta_{n}\eta^{2}}{b}\alpha^{2}$. The recursive formula equation 48 implies that $c_{k}\leq c_{0}$ for any $k$, and
@@
where we have used the fact that $\delta_{n}\leq 1$, Substituting equation 50 into equation 49, we have
@@
We recall from equation 41 that
@@
From equation 51 and the definition of $\beta$, we have
@@
Substituting equation LABEL:eq:gamma_first_bound and equation 55 into equation 53, we obtain
@@
Next, we find the upper bound on $\chi_{m}$ in equation 41 given the current parameter setting and $c_{k}\leq c_{0}$,
@@
where in the big $O$ notation, we only keep the dominant terms and ignore the constant numbers that are independent of $d$, $b$, and $T$ .
Substituting equation 57 and equation 60 into equation 7, we have
@@ Appendix G Proof of Theorem 2
Following together with equation 18, we can obtain that
@@
Substituting equation 63 and equation 30 into equation 62, we have:
@@
Substituting equation 64 into equation 38, we have:
$R_{k+1}^{s}\leq$ $\mathbb{E}\left[f(\mathbf{x}_{k}^{s})\right]\bm{-}\left(\eta_{k}-\frac{c_{k\bm{+}1}\eta_{k}}{\beta_{k}}\right)\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]+\left(c_{k\bm{+}1}+c_{k\bm{+}1}\beta_{k}\eta_{k}\right)\mathbb{E}\left[\|\mathbf{x}_{k}^{s}-\mathbf{x}_{0}^{s}\|_{2}^{2}\right]$ (65)
@@
Based on the definition of $c_{k}=\left(1+\beta_{k}\eta_{k}\right)c_{k\bm{+}1}+\left(\frac{L}{2}+c_{k+1}\right)\left(1+\frac{4d}{q}\right)\frac{12L^{2}\delta_{n}\eta_{k}^{2}\alpha^{2}}{b}$ and $R_{k}^{s}$ given by equation 37, we can simplify equation 65 to
$R_{k+1}^{s}$ $\leq R_{k}^{s}\bm{-}\left(\eta_{k}-\frac{c_{k\bm{+}1}\eta_{k}}{\beta_{k}}\right)\mathbb{E}\left[\|\nabla f(\mathbf{x}_{k}^{s})\|_{2}^{2}\right]$ (66)
@@
Based on equation 66 and the following argument in, we can achieve
@@
[…]
Substituting equation 69 into equation 68, we have:
@@
The second inequality holds if we let $\rho\leq\frac{1}{120}$ Substituting equation 73 and equation 72 into equation 71, we can get
@@
Substituting equation 74 and equation 77 into equation 7, we have